# Which Reasoning Trajectories Teach Students to Reason Better? A Simple Metric of Informative Alignment

### v3

@@ 8 Broader Applicability of RSR
Although our work focuses solely on reasoning tasks, RSR is not specifically designed for reasoning trajectories (CoTs) and can be applied to general text data. Exploring its behavior and effectiveness beyond reasoning settings remains an interesting direction for future work.
@@
Beyond trajectory and teacher selection, RSR may also be applicable to subset selection. Subset selection aims to identify a high-quality subset from a dataset composed of different reasoning problems and their corresponding trajectories. Unlike trajectory selection, which compares multiple trajectories for the same problem, subset selection requires cross-problem comparison of RSR values to filter samples across heterogeneous reasoning prompts. This scenario is particularly relevant when only a single trajectory is available per problem, yet improved data efficiency is still desired. Although subset selection introduces additional confounding factors due to variation across problems—making it harder to isolate trajectory–student suitability effects—RSR may still help identify relatively more effective samples across different reasoning problems. Preliminary experiments on internal data provide suggestive evidence for this possibility.
@@
In addition, the derived metric takes a simple and intuitive form with a clear interpretation. However, it remains unclear whether it arises from deeper mathematical principles. We have not yet identified a suitable theoretical framework to analyze this aspect, which we leave for future investigation.
Finally, due to resource constraints, we primarily focus on mathematical reasoning tasks with extensive controlled studies, while also including additional evaluation on GPQA. Extending our evaluation to other domains,such as code or additional forms of reasoning, would be valuable. However, since our analysis relies on large-scale trajectory generation and approximately 200 SFT experiments, such extensions would require substantially greater computational resources and are therefore left for future work.
@@ A.1 Determining Training Problem Set
We fix the training set size at 5,000 problems for three reasons. First, prior studies [45, 29] have shown that strong reasoning capabilities can be learned from training sets of around 1,000 problems. Second, high-quality reasoning problems are relatively scarce, making further scaling often unrealistic in practice. Third, our distillation study incurs quadratic computational costs across teacher–student pairs. Considering these factors, we find 5,000 problems to offer a good trade-off between representativeness and computational feasibility.
@@ A.3 Teacher Models and Student Models
These teachers produce trajectories that vary in length, informational content, and style. For example, trajectories generated by GPT-OSS models tend to be concise, whereas those from DeepSeek-R1 are generally more detailed. Specific trajectory examples from different teachers are available in our released datasets on Hugging Face and in our case study (§ C.5).
@@ A.4 Benchmark Evaluation
We use vLLM together with the Math-Verify package to evaluate post-trained models on mathematics benchmarks. Our evaluated benchmarks, AIME’25, AIME’24, AMC’23, and MATH500, are widely used and span varying difficulty levels, assessing mathematical reasoning and multi-step problem-solving across diverse domains. We additionally conduct evaluation on GPQA-Diamond beyond mathematics, as described in § C.4.2. We adopt the Acc@4 metric as the final score, which averages results over four independent evaluations per problem. Evaluating a single model checkpoint typically takes around half an hour using 8 H200 GPUs.
@@ A.7 Details of Correlation Analysis
We use the Spearman correlation coefficient [47] because our analysis focuses on monotonic consistency rather than strict linear relationships between metrics and post-training performance. We also report Pearson correlation [7] results in § C.2.
@@ A.8 Definition of Dataset-level RSR
We compare the correlation strength of the simple dataset-level average of trajectory-level RSR, i.e., $\frac{1}{|\mathbf{X}|}\sum_{j=1}^{|\mathbf{X}|}\mathrm{RSR}(\mathbf{x}_{j})$, with that of the dataset-level RSR computed via surprisal-weighted averaging (Eq. 9) in the additional ablation study (§ C.3).
@@ B.1 LLM-judged Quality
We use Qwen3-32B-Instruct in a Non-Thinking configuration as an automatic judge to evaluate reasoning trajectories. Under a fixed evaluation prompt, the judge produces a structured assessment at two levels. First, it assigns dimension-wise scores over five criteria—Factual Accuracy, Logical Rigor, Solution Completeness, Reasoning Efficiency, and Presentation Quality—each accompanied by a brief justification. Second, it aggregates the dimensional assessments into an overall score in $[0,1]$ together with a concise rationale. For dataset-level comparison, we report the mean of the trajectory-level overall scores. The complete evaluation prompt is given in Table 28, adapted from [10].
@@ B.6 Others
Appendix C More Results and Analysis
C.1 Computational Cost for Rank-Surprisal Ratio
The computation of Rank-Surprisal Ratio requires only a single forward pass through the student model, during which we collect each token’s surprisal and clipped rank from the model logits. For a single trajectory, computing token-level surprisals and ranks has a worst-case time complexity of $\mathcal{O}(TV)$, where $T$ denotes the number of response tokens and $V$ is the vocabulary size. This computation does not introduce additional complexity beyond that inherent in the dimensionality of the model logits. Moreover, since RSR only depends on clipped ranks, the rank computation can be efficiently implemented via top-$k$ selection rather than full-vocabulary comparisons, reducing practical runtime and GPU memory usage.
In practice, computing RSR over the 5,000-trajectory dataset with a context length of 32,768 using a 7B model typically takes under one hour on a single H200 GPU with FlashAttention-2, which is significantly cheaper than SFT.
C.2 Additional Results for Correlation Analysis
@@
Table 23, 24, 25, 26, and 27 present the full metric assessment results across different teacher trajectory datasets on Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B, respectively.
C.3 Additional Ablation Study
@@
Table 11: Additional ablation study for Rank-Surprisal Ratio. $\Delta$ denotes the change in average correlation.
We conduct some additional ablation studies. Results are shown in Table 11.
@@
The “Avg-Rank (clipped)” variant applies rank clipping with $r_{max}=100$ to Avg-Rank; however, it still exhibits a notable performance gap compared with RSR. This suggests that rank information alone is insufficient to yield strong correlation, underscoring the importance of jointly leveraging rank and surprisal. The “Rank Minus Surprisal” metric is computed by subtracting the surprisal from the (clipped) rank. Although it also models the relationship between rank and surprisal, this simple subtraction does not align well with post-training performance. "Rank–Entropy Ratio" is defined as the ratio between the average (clipped) token rank and the average token entropy, where entropy is loosely related to surprisal. While this metric achieves relatively strong correlation, its performance remains below that of RSR.
@@
C.4 Additional Results for Trajectory Selection
C.4.1 RSR-based Trajectory Selection Variants and No-Selection Baseline
@@
Table 12 also presents trajectory selection results for two selection method variants based on Rank-Surprisal Ratio, offering further insights.
@@
For "Fewer Candidates per Teacher", we evaluate an 11-to-1 setting in which each candidate pool contains 11 teacher trajectories (one per teacher), instead of the original 33-to-1 setting (three per teacher). This setting focuses on selecting the best trajectory for each problem across different teachers, rather than across multiple generations from the same teacher. The results are comparable, with a slight performance gap relative to the 33-to-1 setting, indicating that RSR remains effective even when each teacher provides only a single trajectory. These findings further suggest that while RSR captures suitability differences across generations from the same teacher, such differences are less pronounced than those across different teacher models.
C.4.2 Additional Evaluation on GPQA and Full Results
@@
Table 13: Comparison of different trajectory selection methods on the GPQA-Diamond benchmark across student models.
@@
The complete trajectory selection results underlying Table 6 are presented in Table 18, 19, 20, 21, and 22.
C.4.3 Analysis of Datasets Selected by RSR
@@
Table 14 shows the data composition of datasets selected by RSR across 11 teacher models. The resulting distributions vary across student models, demonstrating the metric’s ability to select different teacher trajectories tailored to different students. QwQ-32B is generally preferred across student models, consistent with its stable performance. For a clearer contrast in data composition among student models, we refer readers to § C.4.4, where trajectory selection is performed with fewer teachers and consistently strong teachers such as QwQ-32B are removed. We release the RSR-selected datasets for five student models on Hugging Face.
The average RSR values of the selected datasets are also reported in Table 14. These datasets consistently achieve substantially lower RSR values than the teacher trajectory datasets (see § C.2 for complete metrics), validating that our selection procedure effectively identifies trajectories with low RSR for each problem.
C.4.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
To better reflect practical scenarios in which only a few teachers’ trajectories are available and generally suitable teachers may be absent, we conduct additional experiments that select trajectories from a reduced set of teachers. Specifically, we select trajectories from candidates generated by seven teachers: DeepSeek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, GPT-OSS-20B, and Qwen-3-8B. This teacher set is formed by combining the teachers used in Table 2 and 7.
@@ C.4.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
Table 16: Data composition of trajectory datasets selected by RSR under a reduced-teacher setting (see the setting in § C.4.4 and results in Table 15).
C.5 What Does RSR Prefer? A Token-Level Case Study
[…]
We also observe that token-level RSR may become numerically unstable when extremely low surprisal yields inflated RSR values (e.g., 100+). This motivates our surprisal-weighted averaging for trajectory-level RSR (§ 3.4), which stabilizes the metric and improves its effectiveness.
[…]
Table 17: A case study with visualization of token-level rank, surprisal, and RSR values (with rank clipping) across trajectories generated by GPT-OSS-20B, Nemotron-Super-49B-v1.5, and QwQ-32B, as measured by the Qwen-2.5-7B student. Each token is color-coded according to its RSR value, where darker red indicates lower values (stronger preference under RSR). GPT-OSS-20B exhibits uncommon phrasing and high rank values; Nemotron-Super exhibits conventional reasoning flow and low surprisal; QwQ-32B strikes a more favorable balance between informativeness and alignment, yielding lower token-level RSR values. Trajectory-level RSR is defined as the surprisal-weighted average of token-level RSR values with rank clipping (§ 3.4).
Appendix D Others
D.1 License for Artifacts and Data Consent
All artifacts used in this paper are publicly available for academic research purposes, including AIME, AMC, MATH500, and NuminaMath.
D.2 Data Statement
@@
D.3 AI Assistant Usage Statement
@@
Table 18: Full post-training evaluation results for trajectory selection on Qwen-3-14B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
@@
Table 19: Full post-training evaluation results for trajectory selection on LLaMA-3.1-8B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
@@
Table 20: Full post-training evaluation results for trajectory selection on Qwen-2.5-7B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
@@
Table 21: Full post-training evaluation results for trajectory selection on Qwen-3-4B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
@@
Table 22: Full post-training evaluation results for trajectory selection on Qwen-2.5-3B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
@@
Table 28: Evaluation prompt for LLM-judged quality assessment.

### v4

@@ 7 Broader Applicability of RSR
Although our work focuses solely on reasoning tasks, RSR is not specifically designed for reasoning trajectories (CoTs) and can be applied to general text data. Exploring its behavior and effectiveness beyond reasoning settings remains an interesting direction for future work, and we welcome efforts from the community in this direction.
@@
Beyond trajectory and teacher selection, RSR may also be applicable to subset selection. Subset selection aims to identify a high-quality subset from a dataset composed of different reasoning problems and their corresponding trajectories. Unlike trajectory selection, which compares multiple trajectories for the same problem, subset selection requires cross-problem comparison of RSR values to filter samples across heterogeneous reasoning prompts. This scenario is particularly relevant when only a single trajectory is available per problem, yet improved data efficiency is still desired. Although subset selection introduces additional confounding factors due to variation across problems—making it harder to isolate trajectory-student suitability effects—RSR may still help identify relatively more effective samples across different reasoning problems. Preliminary experiments on internal data provide suggestive evidence for this possibility.
8 Conclusion
In this paper, we study data-student suitability in reasoning distillation and propose Rank-Surprisal Ratio (RSR), a simple metric for identifying suitable reasoning trajectories for a given student. Motivated by our analysis, RSR jointly captures a trajectory’s informativeness and alignment with the student’s behavior, favoring trajectories with low absolute probability but relatively high-ranked tokens. Experiments across diverse teacher-student pairs show that RSR strongly correlates with post-training performance and consistently outperforms existing metrics. We further demonstrate its effectiveness in both trajectory and teacher selection. Overall, our results highlight informative alignment as a promising direction for reasoning distillation.
@@ 8 Conclusion
In addition, given the simple and interpretable form of the derived RSR metric, it is natural to ask whether it can be grounded in deeper theoretical principles. However, we have not yet identified a suitable theoretical framework to characterize RSR, which we leave for future investigation.
Finally, due to resource constraints, we focus primarily on mathematical reasoning tasks to conduct extensive controlled studies. Although we perform additional evaluation on GPQA to assess the generalizability of RSR, we acknowledge that the current experiments still have limitations. Specifically, we have not yet applied RSR to newly generated trajectories beyond the existing mathematical reasoning data, nor to qualitatively different domains such as code or commonsense reasoning. Systematic cross-domain studies would be valuable, but they require large-scale trajectory regeneration and retraining, entailing substantial computational cost. Therefore, we focus on the current setting (following prior work) and leave such systematic extensions for future work.
@@ A.1 Determining Training Problem Set
We fix the training set size at 5,000 problems for three reasons. First, prior studies [46, 29] have shown that strong reasoning capabilities can be learned from training sets of around 1,000 problems. Second, high-quality reasoning problems are relatively scarce, making further scaling often unrealistic in practice. Third, our distillation study incurs quadratic computational costs across teacher-student pairs. Considering these factors, we find 5,000 problems to offer a good trade-off between representativeness and computational feasibility.
@@ A.3 Teacher Models and Student Models
These teachers produce trajectories that vary in length, informational content, and style. For example, trajectories generated by GPT-OSS models tend to be concise, whereas those from DeepSeek-R1 are generally more detailed. Specific trajectory examples from different teachers are available in our released datasets on Hugging Face and in our case study (§ D.2).
@@ A.4 Benchmark Evaluation
We use vLLM together with the Math-Verify package to evaluate post-trained models on mathematics benchmarks. Our evaluated benchmarks, AIME’25, AIME’24, AMC’23, and MATH500, are widely used and span varying difficulty levels, assessing mathematical reasoning and multi-step problem-solving across diverse domains. We additionally conduct evaluation on GPQA-Diamond beyond mathematics, as described in § E.3.2. We adopt the Acc@4 metric as the final score, which averages results over four independent evaluations per problem. Evaluating a single model checkpoint typically takes around half an hour using 8 H200 GPUs.
@@ A.7 Details of Correlation Analysis
We use the Spearman correlation coefficient [48] because our analysis focuses on monotonic consistency rather than strict linear relationships between metrics and post-training performance. We also report Pearson correlation [7] results in § E.1.
@@ A.8 Definition of Dataset-level RSR
We compare the correlation strength of the simple dataset-level average of trajectory-level RSR, i.e., $\frac{1}{|\mathbf{X}|}\sum_{j=1}^{|\mathbf{X}|}\mathrm{RSR}(\mathbf{x}_{j})$, with that of the dataset-level RSR computed via surprisal-weighted averaging (Eq. 9) in the additional ablation study (§ E.2).
@@ B.1 LLM-judged Quality
We use Qwen-3-32B-Instruct in a Non-Thinking configuration as an automatic judge to evaluate reasoning trajectories. Under a fixed evaluation prompt, the judge produces a structured assessment at two levels. First, it assigns dimension-wise scores over five criteria—Factual Accuracy, Logical Rigor, Solution Completeness, Reasoning Efficiency, and Presentation Quality—each accompanied by a brief justification. Second, it aggregates the dimensional assessments into an overall score in $[0,1]$ together with a concise rationale. For dataset-level comparison, we report the mean of the trajectory-level overall scores. The complete evaluation prompt is given in Table 30, adapted from [10].
@@ B.6 Others
Appendix C Computational Cost of RSR
The computation of Rank-Surprisal Ratio requires only a single forward pass through the student model, during which we collect each token’s surprisal and clipped rank from the model logits. For a single trajectory, computing token-level surprisals and ranks has a worst-case time complexity of $\mathcal{O}(TV)$, where $T$ denotes the number of response tokens and $V$ is the vocabulary size. This computation does not introduce additional complexity beyond that inherent in the dimensionality of the model logits. Moreover, since RSR only depends on clipped ranks, the rank computation can be efficiently implemented via top-$k$ selection rather than full-vocabulary comparisons (see our code on GitHub), reducing practical runtime and GPU memory usage.
In practice, computing RSR over the 5,000-trajectory dataset with a context length of 32,768 using a 7B model typically takes under one hour on a single H200 GPU with FlashAttention-2 enabled, which is significantly cheaper than SFT.
We also conduct a direct wall-clock time comparison for computing RSR, G-Norm, Influence Score, and LLM-judged Quality on a fixed dataset of 5,000 DeepSeek-R1 trajectories using the LLaMA-3.1-8B student model. As shown in Table 10, RSR incurs significantly lower total GPU-hours than the other metrics. This likely stems from the fact that RSR requires only a single forward pass through the student model, whereas the other methods involve gradient computations or text generation.
Metrics #GPU Hours Total GPU-hours
RSR 1 1 1
G-Norm 2 3 6
Influence Score 8 4 32
LLM-judged Quality 8 1.5 12
Table 10: Computational cost comparison of different metrics on a fixed set of 5,000 DeepSeek-R1 trajectories using the LLaMA-3.1-8B student model, evaluated on H200 GPUs.
Appendix D What Does RSR Prefer? Statistical Analysis and Case Study
D.1 Statistical Analysis of RSR-Selected Trajectories
We compare the statistics of the RSR-selected trajectories with the aggregate statistics (mean/max/min) across source teacher datasets. From Table 11, the RSR-selected trajectories exhibit moderate token length and verified accuracy. This suggests that RSR is not driven solely by length or accuracy, but instead follows its own evaluation criterion. Additional analysis of datasets selected by RSR can be found in § E.3.3.
Datasets Avg. Token Length Verified Accuracy
RSR Selected 9845 81.19
Mean over Teachers 8912 81.58
Max. 13145 85.9
Min. 2552 77.6
Table 11: Statistics of RSR-selected trajectories and source teacher trajectories in terms of token length and verified accuracy. Qwen-2.5-7B is used as the student model. RSR follows its own evaluation criterion, independent of length or accuracy.
D.2 Token-Level Case Study
[…]
We also observe that token-level RSR may become numerically unstable when extremely low surprisal yields inflated RSR values (e.g., >100). This motivates our surprisal-weighted averaging for trajectory-level RSR (§ 3.4), which stabilizes the metric and improves its effectiveness.
[…]
Table 12: A case study with visualization of token-level rank, surprisal, and RSR values (with rank clipping) across trajectories generated by GPT-OSS-20B, Nemotron-Super-49B-v1.5, and QwQ-32B, as measured by the Qwen-2.5-7B student. Each token is color-coded according to its RSR value, where darker red indicates lower values (stronger preference under RSR). GPT-OSS-20B exhibits uncommon phrasing and high rank values; Nemotron-Super exhibits conventional reasoning flow and low surprisal; QwQ-32B strikes a more favorable balance between informativeness and alignment, yielding lower token-level RSR values and thus being more likely to be selected. Trajectory-level RSR is defined as the surprisal-weighted average of token-level RSR values with rank clipping (§ 3.4).
Appendix E More Results and Analysis
E.1 Additional Results for Correlation Analysis
@@
E.2 Additional Ablation Study
We conduct additional ablation studies with more design variants of RSR. Results are shown in Table 14.
@@ E.2 Additional Ablation Study
Avg-Entropy $0.495$ 0.361-0.361
Surprisal-Weighted Sum of Rank $0.604$ 0.252-0.252
Rank-Weighted Sum of Surprisal $0.584$ 0.272-0.272
@@
Table 14: Additional ablation study for Rank-Surprisal Ratio. $\Delta$ denotes the change in average correlation. All ranks are clipped at $r_{max}=100$.
@@
The “Avg-Rank (clipped)” variant applies rank clipping with $r_{max}=100$ to Avg-Rank (see Table 4); however, it still exhibits a notable performance gap compared with RSR. This suggests that rank information alone is insufficient to yield strong correlation, underscoring the importance of jointly leveraging rank and surprisal. Similarly, other metrics, such as average entropy and weighted-sum combinations, yield substantially lower correlation than RSR. Our analysis suggests that these alternatives still tend to emphasize high-likelihood trajectories while insufficiently capturing informativeness.
The “Rank Minus Surprisal” metric is computed by subtracting surprisal from the (clipped) rank. Although it models the relationship between rank and surprisal, this simple subtraction does not align well with post-training performance; similarly, “Rank Times Surprisal” also shows poor alignment. The “Rank-Entropy Ratio” is defined as the ratio between the average (clipped) token rank and the average token entropy, where entropy is loosely related to surprisal. Notably, it exhibits a smaller performance drop, suggesting that ratio-based formulations better balance alignment and informativeness.
@@
E.3 Additional Results for Trajectory Selection
E.3.1 RSR-based Trajectory Selection Variants and No-Selection Baseline
@@
Table 15 also presents trajectory selection results for two selection method variants based on RSR, offering further insights.
@@
For "Fewer Candidates per Teacher", we evaluate an 11-to-1 setting in which each candidate pool contains 11 teacher trajectories (one per teacher), instead of the original 33-to-1 setting (three per teacher). This setting focuses on selecting the best trajectory for each problem across different teachers, rather than across multiple generations from the same teacher. The results are comparable, with a slight performance gap relative to the 33-to-1 setting, indicating that RSR remains effective even when each teacher provides only a single trajectory. The results also suggest that while RSR captures suitability differences across generations from the same teacher, such differences are less pronounced than those across different teacher models.
E.3.2 Additional Evaluation on GPQA
@@
Table 16: Comparison of different trajectory selection methods on the GPQA-Diamond benchmark across student models. We report Acc@4 as the evaluation metric.
@@
Complete tables combining mathematical and GPQA evaluation results can be found in § F.
E.3.3 Analysis of Datasets Selected by RSR
Table 17 shows the data composition of datasets selected by RSR across 11 teacher models. The resulting distributions vary across student models, demonstrating the metric’s ability to select different teacher trajectories tailored to different students. QwQ-32B is generally preferred across student models, consistent with its stable performance. For a clearer contrast in data composition among student models, we refer readers to § E.3.4, where trajectory selection is performed with fewer teachers and consistently strong teachers such as QwQ-32B are removed. We release the RSR-selected datasets for five student models on Hugging Face.
The average RSR values of the selected datasets are also reported in Table 17. These datasets consistently achieve substantially lower RSR values than the teacher trajectory datasets (see § F), validating that our selection procedure effectively identifies trajectories with low RSR for each problem.
@@ E.3.3 Analysis of Datasets Selected by RSR
E.3.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
To better reflect practical scenarios in which only a few teachers’ trajectories are available and generally suitable teachers may be absent, we conduct additional experiments that select trajectories from a reduced set of teachers. Specifically, we select trajectories from candidates generated by seven teachers: DeepSeek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, GPT-OSS-20B, and Qwen-3-8B. This teacher set is formed by combining the representative teachers used in Table 2 and § 5.2.
@@ E.3.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
Table 19: Data composition of trajectory datasets selected by RSR under a reduced-teacher setting (see the setting in § E.3.4 and results in Table 18).
Appendix F Complete Results Tables
Table 25, 26, 27, 28, and 29 present the full metric assessment results across different teacher trajectory datasets on Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B, respectively.
The complete trajectory selection evaluation results underlying Table 6 are presented in Table 20, 21, 22, 23, and 24.
Appendix G Others
G.1 License for Artifacts and Data Consent
All artifacts used in this paper are publicly available for academic research purposes, including AIME, AMC, MATH500, GPQA-Diamond and NuminaMath.
G.2 Data Statement
@@
G.3 AI Assistant Usage Statement
@@
Table 20: Full post-training reasoning evaluation results for trajectory selection on Qwen-3-14B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
@@
Table 21: Full post-training reasoning evaluation results for trajectory selection on LLaMA-3.1-8B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
@@
Table 22: Full post-training reasoning evaluation results for trajectory selection on Qwen-2.5-7B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
@@
Table 23: Full post-training reasoning evaluation results for trajectory selection on Qwen-3-4B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
@@
Table 24: Full post-training reasoning evaluation results for trajectory selection on Qwen-2.5-3B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
@@
Table 30: Evaluation prompt for LLM-judged quality assessment as a baseline metric.